OpenStudy (anonymous):

Find the value of x: 2x^3 - 5x^2 - x + 20 = 0

OpenStudy (anonymous):

would you like to go for a trignometric solution ?

OpenStudy (anonymous):

I just need the value of x :o the answers have 2 x values. Its a whole different equation but Ive added everything and I just needed help on finding the x values of this because It confused me

OpenStudy (anonymous):

what is the ans.?

OpenStudy (anonymous):

if you want soln,,i can provide hint : see shubhamsrg's soln : http://openstudy.com/users/shubhamsrg#/updates/4fed5f8ae4b0bbec5cfcc1ac

OpenStudy (anonymous):

first you have to convert the eqn into depressed but..

OpenStudy (anonymous):

1: x= -4/3, x= -5 2: x= -4/3, x= 5 3:x= 4/3, x= -5 4:x= 4/3, x= 5

OpenStudy (anonymous):

I tried plugging it into wolframalpha already.. didnt work -_-

OpenStudy (anonymous):

didn't work! there the ans. given is -1.6

OpenStudy (anonymous):

Ok that gives me one answer but :o

OpenStudy (anonymous):

I guess ill just try Plugging 5 and negative 5 into the original thing.. Thanks :p

OpenStudy (anonymous):

Wait can Someone actually Help me? Cause Ive tried plugging the two 5's in but Its not equaling out. @Wired @JohnHanShanghai , the original equation is this,

OpenStudy (anonymous):